A particle P of mass 1.2kg is released from rest at the top of a slope and start to move.The slope as length 4m and is inclined at 25degree to the horizontal.The coefficient of friction between P and the slope is 0.25. find (i) the frictional component of the contact force on p (ii) the acceleration (iii) the speed with which P reaches the buttom of the slope.( Take g=10m/s)
1
Expert's answer
2019-10-10T10:01:07-0400

(i) The frictional component of the contact force on P


"F_{\rm fr}=\mu N=\mu mg\cos\theta"

"=0.25\times 1.2\times 10\times \cos 25^{\circ}=2.72\: \rm N"

(ii) The acceleration


"a=\frac{mg\sin\theta-F_{\rm fr}}{m}"

"=\frac{1.2\times 10\times\sin 25^{\circ}-2.72}{1.2}=1.96\:\rm m/s^2"

(iii) The speed with which P reaches the bottom of the slope


"v=\sqrt{2ad}=\sqrt{2\times 1.96\times 4}=3.96\:\rm m/s"